OpenStudy (anonymous):

-2(5x+2)-3(-3x-5)=10

OpenStudy (anonymous):

Okay, first distribute the -2 to 5x+2, and -3 to -3x-5. Combine like terms. Then Isolate x.

OpenStudy (anonymous):

-10x-4+9x+15, -x+11=10 -x=-1 x=1